Consider the battery charging circuit with Vm = 20 V, R = 10 ohms and VB = 14 V. Find the peak current assuming an ideal diode. Also, find the percentage of each cycle in which the diode is in on state. Sketch Vs(t) and i(t) to scale against time.


1
Expert's answer
2021-10-10T16:05:05-0400

The diode is on when

"V_msin(\omega t)>V_B" or "20sin(\omega t)>14"

This diode goes to on state at

"20sin(\omega t)=14"

"\omega t=arcsin(14/20)=arcsin0.7"

"\omega t=45\degree;135\degree"

The diode is on for "45\degree\le\omega t\le135\degree" or for "90\degree" of the phase angle. The whole period is "360\degree", so the diode is on for

"90\degree/360\degree=0.25=25\%" of the time.

The peak current is when the ac voltage is at the peak and is

"I_m=\frac{V_m-V_B}{R}=\frac{20-14}{10}=0.6" A